### v3

# AutoSAM: Towards Automatic Sampling of User Behaviors for Sequential Recommender Systems

Abstract.
Sequential recommender systems (SRS) have garnered immense popularity in recommendation due to their remarkable proficiency in capturing dynamic user preferences. In the current setup of SRS, a common configuration is to uniformly consider each historical behavior as a positive interaction. However, this setting has the potential to yield sub-optimal performance as each individual item often have a different impact on shaping the user’s interests. For instance, purchased items should hold more significance than merely clicked ones. Hence, in this paper, we propose a novel automatic sampling framework for sequential recommendation, named AutoSAM, to non-uniformly treat historical behaviors. Specifically, AutoSAM extends the conventional SRS framework by integrating an extra sampler to intelligently discern the skew distribution of the raw input, and then sample informative sub-sets to build more generalizable SRS. To tackle the challenges posed by non-differentiable sampling actions and to introduce multiple decision factors for sampling, we further design a novel reinforcement learning based method to guide the training of the sampler. We theoretically devise multi-objective sampling rewards including Future Prediction and Sequence Perplexity, and then optimize the whole framework in an end-to-end manner by combining the policy gradient. We conduct a broad range of tests on benchmark recommendation models and four real-world datasets from different online platforms. The experimental results demonstrate the effectiveness of the proposed approach. We also make our code publicly available  [^1] https://github.com/zh-ustc/AutoSAM.
Sequential Recommendation, User Interest, Automatic Sampling
Qi Liu is corresponding author.
conference: XXX; XXX; ccs: Information systems Recommender System
1. Introduction
Recommender systems  (Koren, 2008; Koren et al., 2009; Liu et al., 2018; Wu et al., 2022) have become crucial tools for information filtering in various online applications, such as e-commerce, advertising, and online videos. Among them, sequential recommender systems (SRS)  (Huang et al., 2018; Wang et al., 2019; Li et al., 2017; Liu et al., 2022; Singer et al., 2022) have become increasingly prevalent due to their ability to capture long-term and short-term user interests.
So far, many efforts have been devoted to sequential recommendation, ranging from early matrix factorization and Markov chain based method (Rendle et al., 2010) to state-of-the-art deep neural network models, including recurrent neural networks (RNNs) (Hidasi et al., 2015, 2016), convolutional neural networks (CNNs) (Tang and Wang, 2018; Yuan et al., 2019), and self-attentive models (Kang and McAuley, 2018; Sun et al., 2019). Meanwhile, the effectiveness and efficiency of the generative loss (i.e., auto-regressive loss) (Yuan et al., 2019) in SRS tasks have been widely demonstrated in these approaches. Despite their remarkable success, these methods often consider from a model perspective while treating all historical behaviors as uniformly positive. In fact, this may lead to sub-optimal performance as items in the sequences are usually of unequal importance on shaping user interests (Zhang et al., 2013). That is because that the motivation behind the user behaviors are often very complex and multifaceted. For example, some interactions on online platforms may be driven by a herd mentality, specific purchasing needs or even the result of mis-clicks, while others may represent the user’s true preferences.
Recently, there have also been some works which attempt to improve the recommender systems from a data perspective (He et al., 2023; Qin et al., 2023). To be specific, SIM (Pi et al., 2020) extracts user interests with two cascaded search units to capture the diverse user’s long-term interest with target item. Similarly, UBR4CTR (Qin et al., 2020) adopts two-stage frameworks to model long-term interests. At the first stage, it generates the query according to the target item for retrieving similar historical behaviors. Then, at the second stage, these items are leveraged to predict the user’s interests. To reduce the information loss of the above methods, SDIM (Cao et al., 2022) proposes a sampling-based end-to-end approach for CTR prediction, which samples from multiple hash functions to gather similar behavior items to the target. Though the aforementioned methods have proven effective, they may suffer from the following limitations. Firstly, these methods often select items according to the target, which brings challenges to compute scores across all candidate items in parallel and may be only suitable to fine-ranking stage (Covington et al., 2016). Besides, such a setup may be too strict and fail to consider from a dynamic perspective. Secondly, these sampling processes are mainly designed for CTR prediction tasks (Guo et al., 2017; Wang et al., 2017), which are often typically optimized as binary classification problems. As a result, these approaches often cannot be equipped with generative loss in SRS tasks, whose effectiveness and efficiency have been demonstrated in many previous works (Kang and McAuley, 2018; Yuan et al., 2019).
We hold that an ideal solution should be adaptive while maintaining the core training principles of SRS. Hence, to achieve this goal, we propose a general automatic sampling framework, named AutoSAM, to non-uniformly treat historical behaviors for sequential recommendation. To be concrete, an additional sampler layer is first employed to adaptively explore the skew distribution of raw input. Then, as shown in Figure 1, informative sub-sets are sampled from the distribution to build more generalizable sequential recommenders. Meanwhile, the removal of some low-signal items releases the model from the constraint of always taking the next step as the target in traditional SRS, which helps explore more potential preferred items in the future.
Figure 1. An illustration of a user’s generated samples, items outlined by red boxes stand for the user’s favorite, while the green box means misclicks, etc.
In order to overcome the challenges of non-differentiable discrete sampling actions, and to introduce multiple decision factors for sampling, we further introduce a novel reinforcement learning based method to guide the training of the sampler. We believe that a proper decision should not only focus on the target item in the future, but also consider the coherence with the previous context. Along this line, we incorporate the major factors into the reward estimation, including Future Prediction and Sequence Perplexity. Finally, both the sequential recommender and the sampler can be jointly optimized in an end-to-end manner by combining the policy gradient. As a result, our method exhibits more accurate recommendations than previous approaches, and is generally effective for various sequential recommenders with different backbones. We summarize the contributions as follows:
  • $\bullet$
    We propose to sample user behaviors from a non-uniform distribution for sequential recommendation. We highlight that the main challenge is to sample historical behaviors dynamically while leveraging the benefits of generative loss.
  • $\bullet$
    We propose a general automatic sampling framework for sequential recommendation, named as AutoSAM, to build generalizable SRS with an additional sample layer. We further introduce a reinforcement learning based method to solve the challenges of non-differentiable decisions and design multi-objective rewards to guide the training of the sampler.
  • $\bullet$
    We conduct extensive experiments on public datasets to show superior recommendation results compared to previous competitive baselines. We also validate the generality of the proposed method and additionally analyze the effectiveness of the sampler.
Figure 2. (a) is the architecture of the standard sequential recommender system (SRS). (b) is our proposed AutoSAM which augments the traditional sequential recommendation architecture with an additional sampler layer to adaptively explore the skew distribution of raw input. The dashed lines represent gradient backpropagation.
2. Preliminaries
Before going into the details, we first introduce the research problem and revisit the sequential recommender systems.
2.1. Problem Definition
Assume that there are item set =$\mathcal{I}=\{i_{1},i_{2},...,i_{|\mathcal{I}|}\}$ and user set $\mathcal{U}=\{u_{1},u_{2},...,u_{|\mathcal{U}|}\}$, we denote each behavior sequence of user $u\in\mathcal{U}$ as $X^{u}=[x_{1}^{u},x^{u}_{2},...,x_{n}^{u}]$, where $x_{t}^{u}\in\mathcal{I}$ is the item that $u$ interacted at step $t$ and $n$ is the sequence length. Sequential recommender systems (SRS) aim to predict the item that users might interact with at the next time step. Different from traditional SRS, the main idea behind this work is to sample historical behaviors with a policy $\pi(A^{u}|X^{u})$, where $A^{u}\in\{select=1,discard=0\}^{n}$, and then utilize $\hat{X^{u}}=[x_{k_{1}}^{u},x_{k_{2}}^{u},...,x_{k_{m}}^{u}]$ to train more generalizable SRS, in which $k$ is the set of time steps about sampled items, $m$ is the length of $\hat{X^{u}}$.
2.2. Sequential Recommender System
As shown in Figure 2(a), modern SRS mainly contain the components as following. First, each input item $x^{u}_{t}$ is mapped into an embedding vector by an embedding lookup operation as:
(1) $E_{t}=\zeta_{x_{t}},$
where $\zeta\in\mathbb{R}^{|\mathcal{I}|\times d}$ is item embedding matrix. Then, the sequential embedding will be fed into stacked hidden layers in order to capture the long- and short-term dependence.
(2) $H^{(l)}$ $=F^{(l)}(H^{(l-1)}),1\leq l\leq L,$
where $L$ stands for the number of layers, $H^{(l)}$ is the output of the $l$-th hidden layer $F^{(l)}$ while $H^{(0)}=E$. Next, a prediction layer is adopted to generate the distribution $P$ of the user’s next preferences for each time step $t$, which can be written as:
(3) $P_{t}=Softmax(H^{(L)}_{t}W+b),$
in which $W\in\mathbb{R}^{d\times|\mathcal{I}|}$, $b\in\mathbb{R}^{|\mathcal{I}|}$ is the learnable parameters. Finally, to train the model, the auto-regressive generative loss are employed as optimization objectives, i.e., left-to-right supervision signals. Combining the cross-entropy loss with such an objective, the loss function can be written as follows:
(4) $\mathcal{L}(X;\theta)=-\sum_{t=1}^{n-1}logP(y_{t}|X_{\leq t};\theta)=-\sum_{t=1}^{n-1}logP_{t,y_{t}},$
where $\theta$ denotes the parameters of the model, $y_{t}=x_{t+1}$ is the item expect to be predicted at step $t$, and $n$ is the input length.
3. AutoSAM: the Proposed Method
3.1. Overview of the Framework
The framework of our proposed AutoSAM is depicted in Figure 2(b). First, we employ a light-weighted sampler to adaptively learn the non-uniform distribution of the raw input, and then sample the informative sub-sets from the whole sequence. After that, the sequential recommender systems (SRS) are trained with these higher-quality and diverse samples to gain stronger generalizations. Considering the challenges of non-differentiable sampling actions and the necessity to introduce multiple decision factors, we further introduce a novel reinforcement learning based method to optimize the sampler. Specifically, we treat the sampler as the agent. For each time step $t$ of the user behavior sequence $X$, the sampler takes $X_{\leq t}$ as the current state $s_{t}$, and outputs an action $a_{t}\in\{0,1\}$ of the $t$-th item. Then, it observes the carefully designed multi-objective rewards from the environment to update the model via the policy gradient. Finally, the sampler could make optimal decisions by continuously interacting with the environment. And it is worth mentioning that both the sampler and the SRS can be jointly optimized in an end-to-end manner in our proposed framework.
3.2. User Behavior Sampler
We first transform sequence $X$ into embedding $E_{X}=\zeta_{X}+P_{X}$, where $\zeta$ and $P$ are the embedding matrix and position embedding, respectively. We suggest that the sample decisions should be based on both local and global information. Hence, the sampler leverages a Transformer Block with triangular mask matrix to aggregate the global information $X_{\leq t}$ at each step $t$, and then concatenate it with the local item embedding to generate the sample policy $\pi$ through an MLP layer. The above procedure can be formulated as:
(5) $H$ $=Transformer(E)$
$S$ $=ReLU([H||E]W_{1}+b_{1})W_{2}+b_{2},$
where ||||| | means concatenation, $W_{1}\in\mathbb{R}^{2d\times d},b_{1}\in\mathbb{R}^{d},W_{2}\in\mathbb{R}^{d\times 2},b_{2}\in\mathbb{R}^{2}$ are learnable parameters. After that, binary decisions $A\in\{0,1\}^{n}$ are sampled from the Bernoulli distribution as following:
(6) $\pi=Softmax(S/\tau)\in\mathbb{R}^{n\times 2}\;\;\;\;\;\;\;\;\;\;$
$A\sim\pi\in\{0,1\}^{n},$
where $\tau$ is the temperature of the Softmax. Finally, we can build the sampled sub-set as $\hat{X}=\{x_{t}|a_{t}=1\}$.
3.3. Multi-Objective Rewards
Now we discuss how to design our reward, which plays an important role in learning the optimal $\pi$. As mentioned before, traditional SRS often follow the next item prediction task, which can neither update sampler network directly due to the non-differentiable challenge of discrete actions nor control the sample rate. In addition, such single objective may be too strict and one-sided. We argue that the proper decisions should not only focus on the target item but also consider the coherence with preceding context. Along this line, in this paper, we incorporate the major factors into the reward estimation, including Future Prediction and Sequence Perplexity.
3.3.1. Future Prediction
Generally, items that contribute to predicting the future interactions should be of more opportunities to be selected. Accordingly, we theoretically propose our rewards with consideration of the recommendation loss in Eq. 4. Since we train the SRS with the sampled sub-set $\hat{X}$, the objective can be written as:
(7) $min\;J(\theta,\phi)$ $=min\;\mathbb{E}_{X}\mathbb{E}_{A\sim\pi_{\phi}(A|X)}\left[\mathcal{L}(\hat{X};\theta)\right],$
where $\theta$ and $\phi$ are the parameters of SRS and sampler, respectively, while sharing the embedding layer. To optimize this objective, we first transform the prediction loss to an equivalent form:
(8) $\mathcal{L}(\hat{X};\theta)$ $=\sum_{t=1}^{m-1}-logP(\hat{y}_{t}|\hat{X}_{\leq t};\theta)$
$=\sum_{t=1}^{n-1}-a_{t+1}logP(y_{t}|\{x_{k_{j}}|1\leq j\leq m-1;k_{j}\leq t\};\theta)$
$\overset{\triangle}{=}\sum_{t=1}^{n-1}\mathcal{L}_{t}(X,A;\theta),$
where $m,n$ is the length of $\hat{X},X$ respectively. And then derive the gradients with respect to $\phi$, we have:
(9) $\nabla_{\phi}J(\theta,\phi)$ $=\mathbb{E}_{X}\nabla_{\phi}\sum_{A}\pi_{\phi}(A|X)\mathcal{L}(\hat{X};\theta)$
$=\mathbb{E}_{X}\mathbb{E}_{A\sim\pi_{\phi}(A|X)}\nabla_{\phi}\sum_{t=1}^{n}[log\pi_{\phi}(a_{t}|X)\sum_{t=1}^{n-1}\mathcal{L}_{t}(X,A;\theta)]$
$=\mathbb{E}_{X}\mathbb{E}_{A\sim\pi_{\phi}(A|X)}\nabla_{\phi}\sum_{t=1}^{n}-log\pi_{\phi}(a_{t}|X)r^{pre},$
where $r^{pre}=\sum_{t=1}^{n-1}-\mathcal{L}_{t}(X,A;\theta)$ means the rewards associated to the sampler. More details of mathematical derivations can be found in Appendix A.1. In practice, we may make a few changes to it. First, we add a baseline $\mathcal{L}^{b}$ computed on raw input to $\mathcal{L}$ as it is always positive. Second, we discount the future rewards by a factor $\gamma\in[0,1]$ and ignore the targets whose sample probabilities are lower than $\psi$ to increase stability. Then, the future prediction reward at each time step $t$ can be defined as follows:
(10) $r^{pre}_{t}=\sum_{t^{\prime}=t}^{n-1}-\gamma^{t^{\prime}-t}I[\pi_{\phi}(a_{t^{\prime}+1}|X)>\psi](\mathcal{L}_{t^{\prime}}(X,A;\theta)-\mathcal{L}_{t^{\prime}}^{b}),$
in which $I[\cdot]$ is the indicator function, $\gamma$ is set to $0.9$ across all the experiments in this work.
3.3.2. Sequence Perplexity
In addition to sampling based on future predictions, it should also be attached great importance to considering the coherence with previous context. Inspired by perplexity (PP), which is widely used to evaluate the quality of sentences in Natural Language Processing (NLP) (Liu et al., 2016; Dathathri et al., 2019) with the definition as follows:
(11) $PP(X)=P(x_{1},x_{2},...x_{n})^{-\frac{1}{n}}=\prod_{t=1}^{n}P(x_{t}|X_{<t})^{-\frac{1}{n}}.$
As words causing high perplexity tend to stray from the context, similarly, we insist that the items causing high perplexity may lack representativeness. A natural idea is to assign lower sampling probabilities to the items whose predicted probabilities according to historical behaviors are lower than $1/PP(X)$, so as to reduce the perplexity and obtain informative sub-sets in some way. To this end, the sampler should be encouraged to drop $x_{t}$ ($a_{t}=0$) while $P(x_{t}|X_{<t})<1/PP(X)$, and select $x_{t}$ while $P(x_{t}|X_{<t})>1/PP(X)$.
Note that the perplexity can be transformed to cross-entropy, in practice, we employ the SRS to compute the approximate objective at each time step $t$. Besides, a relax factor is further conducted to control the strictness of the sampler. Above all, the reward can be defined as follow:
(12) $r^{pp}_{t}=\left(logP(x_{t}|X_{<t})-log\frac{1}{PP(X)}+b\right)(2a_{t}-1),$
where $b$ is the relax factor, by which we can indirectly control the global sampling ratio. Generally, the smaller $b$ will lead the sampler to be stricter.
3.4. Optimization
We summarize the rewards with trade-off parameter $\lambda$ and scaling factors $k$ as:
(13) $r=k(\lambda r^{pre}+(1-\lambda)r^{pp}).$
To further learn the sequential recommendation task, we also derive the gradients according to Eq. 7 with respect to $\theta$:
(14) $\nabla_{\theta}J(\theta,\phi)=\mathbb{E}_{X}\mathbb{E}_{A\sim\pi_{\phi}(A|X)}\nabla_{\theta}\mathcal{L}(\hat{X};\theta).$
Finally, both the SRS and sampler can be optimized jointly in an end-to-end manner by combining the policy gradient:
(15) $\theta$ $\leftarrow\theta-\alpha_{1}\nabla_{\theta}\mathcal{L}(\hat{X};\theta)$
$\phi$ $\leftarrow\phi+\alpha_{2}\nabla_{\phi}\sum_{t=1}^{n}log\pi_{\phi}(a_{t}|X)r_{t},$
where $\alpha_{1}$, $\alpha_{2}$ are the learning rates of SRS and sampler respectively. Overall, we present our learning algorithm in Algorithm 1.
3.5. Time and Computation Complexity Analysis
To evaluate the efficiency of our proposed method for online services, we analyze the time complexity of AutoSAM during inference in this part. Similar to the training stage, we first sample representative historical behaviors, and then feed them to SRS for modeling the user sequence. Denote $L$ as the number of layers and $N$ as the sequence length. Since we conduct Transformer-based SRS whose time complexity is $\mathcal{O}(LN^{2})$, the sampling processing takes $\mathcal{O}(N^{2})$ and the sequence modeling can be done within $\mathcal{O}(L(\mathbb{E}[(\sum_{i=1}^{N}a_{i})^{2}]))=\mathcal{O}(L(\mu^{2}N^{2}+N\sigma^{2}))$, in which $\mu\in\{0,1\}$ denotes the average sampling rate while $\sigma$ is the variance. Thus, AutoSAM could reduce the time complexity of about $\mathcal{O}((L-\mu^{2}L-1)N^{2})$ by shorten the behavior sequence. Moreover, since sampling can be regarded as a binary classification problem, we believe the model could be more efficient while maintaining comparable performance by adopting much lighter sampler like RNN or MLP. We will explore such effects in the future work.
3.6. Discussion
Since the existing models like SASRec which leverage attention mechanisms can also adaptive learn the weights of interacted items, we mainly conduct in-depth discussion of the proposed AutoSAM’s advantages compared to these attention-based approaches as the following summary:
  • Firstly, as the auto-regressive loss uniformly takes each item as the supervised signal of the previous sequence, existing attention based methods may still make the low-signal items participate in the training process to a great extent, potentially hurting the model performance. While AutoSAM directly consider the problem from a data perspective, it effectively reduces such risk.
  • Secondly, AutoSAM leverages a reinforcement learning based method to guide the training of the sampler, bringing great flexibility in simultaneously modeling multiple sampling factors. Apart from the rewards deviced in this paper, we believe the proposed framework can also incorporate extra information (e.g. ratings, comments) to design more holistic reward functions.
  • Lastly, by employing hard sampling, AutoSAM can balance effectiveness and efficiency by controlling the sampling rates, which is more conducive to practice.
Algorithm 1 Learning the AutoSAM framework
1:  Initial SRS parameters $\theta$ and sampler parameters $\phi$.
2:  for $Epoch=1,2,...T$ do
3:     for $u=1,2,...|U|$ do
4:        $X\leftarrow[x_{1}^{u},x_{2}^{u},...,x_{n}^{u}]$;
5:        $\hat{X}\leftarrow[\;]$;
6:        Compute the policy $\pi_{\phi}(A|X)\in[0,1]^{n\times 2}$ by sampler;
7:        for $t=1,2,...n$ do
8:           Sample $a_{t}$ from $\pi_{t}$;
9:           if $a_{t}=1$ then
10:              Append $x_{t}$ to $\hat{X}$;
11:           end if
12:        end for
13:        Compute the recommendation loss $\mathcal{L}(\hat{X};\theta)$ according to Eq.4;
14:        Compute overall reward $r$ according to Eq. 10, 12, 13;
15:        Update $\theta$ and $\phi$ according to Eq. 15;
16:     end for
17:  end for
4. Experiments
4.1. Experimental Setup
4.1.1. Datasets
We conduct four real-world datasets from different online platforms as following:
Table 1. Statics of the used datasets in the experiments.
 Datasets #Num. Users #Num. Items #Num. Actions #Actions/Item #Actions/User
Tmall 424,170 969,426 22,010,938 22.71 51.89
Alipay 520,064 2,076,041 20,976,085 10.10 40.33
Yelp 221,397 147,376 3,523,285 23.91 15.91
Amazon 724,012 1,822,885 18,216,875 9.99 25.16
Similar to some previous works (Zhou et al., 2022), we filter inactive users and items with fewer than $c$ interaction records, where $c$ is set to 10 for Tmall, Alipay and Amazon while adjusted to 5 for Yelp due to the smaller data scale. The statics are summarized in Table 1.
Table 2. Recommendation performance comparison of different models over four datasets. The best performance and the second best performance methods are indicated by bold and underlined fonts. “IMP” denotes the improvements of our method compared to the best baseline.
Method Tmall Alipay Yelp Amazon
N@10 N@20 R@10 R@20 N@10 N@20 R@10 R@20 N@10 N@20 R@10 R@20 N@10 N@20 R@10 R@20
PopRec 0.0045 0.0053 0.0087 0.0118 0.0136 0.0143 0.0155 0.0182 0.0037 0.0048 0.0066 0.0111 0.0016 0.0018 0.0030 0.0038
BPR-MF 0.0217 0.0268 0.0402 0.0602 0.0129 0.0142 0.0181 0.0234 0.0173 0.0228 0.0343 0.0564 0.0116 0.0145 0.0212 0.0324
FPMC 0.0379 0.0452 0.0683 0.0971 0.0572 0.0632 0.0907 0.1144 0.0210 0.0273 0.0407 0.0657 0.0481 0.0520 0.0662 0.0817
GRU4Rec 0.0410 0.0503 0.0765 0.1135 0.0450 0.0518 0.0779 0.1050 0.0213 0.0276 0.0415 0.0666 0.0370 0.0421 0.0581 0.0783
NextitNet 0.0425 0.0510 0.0775 0.1115 0.0498 0.0568 0.0850 0.1124 0.0229 0.0295 0.0447 0.0711 0.0441 0.0493 0.0666 0.0863
BERT4Rec 0.0496 0.0595 0.0906 0.1301 0.0485 0.0551 0.0815 0.1078 0.0235 0.0305 0.0460 0.0739 0.0386 0.0441 0.0618 0.0840
SR-GNN 0.0359 0.0437 0.0664 0.0975 0.0401 0.0459 0.0681 0.0915 0.0199 0.0255 0.0383 0.0608 0.0312 0.0355 0.0485 0.0655
CL4SRec 0.0516 0.0616 0.0929 0.1322 0.0616 0.0694 0.1040 0.1351 0.0256 0.0322 0.0486 0.0766 0.0472 0.0528 0.0726 0.0950
DuoRec 0.0526 0.0627 0.0950 0.1350 0.0619 0.0698 0.1050 0.1365 0.0248 0.0317 0.0476 0.0752 0.0478 0.0534 0.0731 0.0954
FullSAM 0.0515 0.0614 0.0927 0.1320 0.0617 0.0695 0.1037 0.1346 0.0245 0.0314 0.0473 0.0746 0.0470 0.0524 0.0718 0.0932
FMLP-Rec 0.0558 0.0665 0.1007 0.1431 0.0630 0.0709 0.1052 0.1362 0.0263 0.0337 0.0506 0.0798 0.0480 0.0535 0.0729 0.0949
RETR 0.0526 0.0624 0.0942 0.1332 0.0608 0.0688 0.1027 0.1346 0.0250 0.0318 0.0478 0.0751 0.0471 0.0525 0.0711 0.0926
SDIM 0.0527 0.0626 0.0939 0.1333 0.0585 0.0685 0.0991 0.1332 0.0241 0.0308 0.0464 0.0734 0.0464 0.0513 0.0701 0.0898
RanSAM 0.0532 0.0633 0.0961 0.1354 0.0633 0.0713 0.1061 0.1380 0.0246 0.0315 0.0475 0.0749 0.0443 0.0500 0.0685 0.0914
LastSAM 0.0502 0.0600 0.0904 0.1291 0.0599 0.0667 0.0993 0.1297 0.0234 0.0297 0.0445 0.0698 0.0446 0.0514 0.0696 0.0893
PopSAM 0.0522 0.0623 0.0945 0.1347 0.0590 0.0665 0.0995 0.1292 0.0246 0.0316 0.0474 0.0752 0.0442 0.0496 0.0681 0.0893
AutoSAM 0.0604 0.0717 0.1084 0.1528 0.0672 0.0753 0.1117 0.1437 0.0272 0.0347 0.0521 0.0817 0.0549 0.0610 0.0831 0.1074
IMP 8.24% 7.82% 7.65% 6.78% 6.16% 5.61% 5.28% 4.13% 3.42% 2.97% 2.96% 2.38% 14.14% 14.02% 13.68% 12.58%
4.1.2. Compared Methods
We first compare our proposed method from data perspective with following approaches:
  • FullSAM trains normally with full historical behaviors.
  • FMLP-Rec (Zhou et al., 2022) conducts Fourier transform with learnable filters to alleviate the influence of noise.
  • RETR (Yao et al., 2022) builds recommender transformer to sample the user behavior pathway for sequential recommendation.
  • SDIM (Cao et al., 2022) replaces the attention mechanism with multiple hash functions to sample relevant historical items for modeling long-term preference.
  • RanSAM, LastSAM, PopSAM samples random, last or most popular behaviors with a fixed probability respectively.
Note that we adopt SASRec (Kang and McAuley, 2018) as the sequential recommender system for all sample-based methods due to its effectiveness. It is worth mentioning that some other sampling based methods e.g. UBR4CTR, SIM are mainly designed for CTR tasks and typically optimized as a binary classification problem. While they sampling historical behaviors using target item, the sampled user behaviors often depend on what the candidate item is (Cao et al., 2022). As a result, these methods are often hard to compute the scores of a large candidate set as the million-level setting in this paper.
Besides, We also compare the models with other architectures:
  • PopRec is a popularity-based method, in which each user is recommended according to the popularity distribution.
  • BPR-MF (Rendle et al., 2012) is a well-known matrix factorization-based method optimized by Bayesian personalized ranking loss.
  • FPMC (Rendle et al., 2010) combines matrix factorization with the Markov chain to predict the next interaction.
  • GRU4Rec (Hidasi et al., 2015) is a pioneering attempts by employing recurrent neural network (RNNs) for next item recommendation.
  • NextitNet (Yuan et al., 2019) is a CNN-based method to capture long-term and short-term interests.
  • BERT4Rec (Sun et al., 2019) utilizes a bidirectional self-attention network to model user sequence.
  • SR-GNN (Wu et al., 2019) applies GNN with attention network to model each session.
  • CL4SRec (Xie et al., 2022b) utilizes contrastive learning to extract the discrimination information in sequential recommendation.
  • DuoRec (Qiu et al., 2022) is proposed with a contrastive objective serving as the regularization over sequence representations.
(a) Tmall
(b) Yelp
Figure 3. Performance comparison (NDCG@10) using different backbones on Tmall and Yelp dataset.
4.1.3. Hyper-parameter Settings
We set the batch size of 128 with 10, 000 random negative items per batch, and the embedding size is set to 128 for all methods. In all sampling based methods which employ two-layer SASRec as the backbone, we conduct 4 multi-head self-attention and 2 FFN layers, while the hidden size is set to 256. The sample rates of RanSAM, LastSAM and PopSAM are searched from $\{0.5,0.6,0.7,0.8,0.9\}$. We consistently employ Autom as the default optimizer for all recommenders, combined with a learning rate $\alpha_{1}$ of $1\times e^{-3}$. As for our sampler, we conduct SGD with a learning rate $\alpha_{2}$ of $1\times e^{-1}$. The remain parameters of baselines are carefully tuned or follow the default settings suggested by the original work. We use grid search to find the best group of AutoSAM’s hyper-parameters as shown in appendix A.2.1.
4.1.4. Evaluation Metrics
We evaluate the recommendation performance over all candidate items with Recall and Normalized Discounted Cumulative Gain (NDCG). The first one is an evaluation of unranked retrieval sets while the other reflects the order of ranked lists. We consider top-k of overall item set for recommendations where $k\in\{10,20\}$ in our experiments. And we split the dataset into training, validation and testing sets following the leave-one-out strategy described in (Cheng et al., 2022; Zhao et al., 2022).
4.2. Performance Comparison
The result of different models on datasets are shown in Table 2. First of all, we can find that considering from the data perspective instead of treating all items as uniformly positive can benefit the performance. However, SDIM achieves limited improvement compared to FullSAM. A possible reason is that SDIM samples according to the last item maybe too strict to capture the user’s rich interests. Besides, FMLP-Rec achieves much better performance than traditional SRS probably due to its stronger ability to filter noise information. And we surprisingly find that RanSAM outperforms FullSAM, this suggests that sampling can be regarded as a method of data augmentation in some way since different samples of each user are generated in different iteration rounds, which lead to significant enhancement in generalization capability by increasing the diversity of the training sequence. Unfortunately, RanSAM treats each interaction uniformly and may also drop lots of high-quality behaviors, making the improvement limited. Different from these baselines, the proposed AutoSAM adaptively explore the skew distribution of the raw input, and then enhance the sequntial recommender with informative samples. Consequently, our method performs best on all datasets, which gains an average improvement over the best baseline of 7.40%, 7.99% on Recall@10, NDCG@10, and 6.47%, 7.60% on Recall@20, NDCG@20 respectively.
4.3. Generality Analysis
To investigate the general applicability of AutoSAM to other sequential recommendation models, we conduct a performance comparison between models using different sampling or filtering methods on various architectures, including RNN-based GRU4Rec (Hidasi et al., 2015), CNN-based NextitNet (Yuan et al., 2019) and Transformer-based SASRec (Kang and McAuley, 2018). It should be noted that since RETR and SDIM are improvements based on attention mechanisms, they are not universally applicable to other architectures. Figure 3 presents the comparison results in terms of test NDCG@10 on Yelp and Tmall datasets. It is evident that all the sequential recommenders demonstrate the most significant improvements through the integration of automatic sampling, highlighting the generality of our AutoSAM for different backbones.
4.4. Evaluation of Compared Methods w.r.t. Number of Test Items
Traditional SRS often treat the item of the next step as the unique prediction target during training and evaluation, which may limit capturing the user’s rich interests. Actually, behaviors after multiple steps may also be in line with the user’s current preferences (Bacciu et al., 2023). Luckily, such a dilemma can be largely alleviated by the proposed AutoSAM. When $x_{t}$ is selected, the distribution of the training target at step $t$ satisfies that: $p(y_{t}=x_{t+n})=p(a_{t+n}=1)\prod_{i=t+1}^{t+n-1}p(a_{i}=0),$ which releases the model from the constraint of always taking the next step as the target in traditional SRS and adaptively broadens the learning horizon. To verify such effect, we test the robustness of longer prediction by extending the test set with multiple time steps. Specifically, we re-split the datasets using time step where 1-st to ($n$-10)-th for training, ($n$-10+1)-th to ($n$-10+$i$)-th and ($n$-5+1)-th to ($n$-5+$i$)-th for validation and testing respectively, in which $i\in\{1,2,3,4,5\}$ denotes the number of items in test set. Performance w.r.t. the number of test items is shown in Table 3. The improvement over baseline substantially increases when more items in the future are added to the test set. This perceptive observation illustrates that AutoSAM frees the model from the constraints of always targeting the next step as standard practice in traditional SRS, thus enhancing the model’s robustness.
Table 3. Recommendation performance (Recall@10) comparison w.r.t. the number of test steps. “IMP” denotes the improvements of our method compared to the baseline.
Dataset Method Num. of Test Steps
1 2 3 4 5
Tmall SASRec 7.99 7.13 6.52 6.09 5.78
AutoSAM 9.10 8.15 7.51 7.04 6.67
IMP 13.89% 14.31% 15.18% 15.60% 15.40%
Amazon SASRec 8.22 6.18 5.00 4.20 3.57
AutoSAM 9.01 7.04 5.80 4.90 4.18
IMP 9.61% 13.92% 16.00% 16.67% 17.09%
Table 4. Ablation study of each reward component, scaling factors $k$ and trade-off parameter $\lambda$ on Alipay dataset. Note that ”w/o r” means using the random initialized sampler throughout. We fix $\lambda$ as $0.5$ when tuning $k$, and fix $k$ as $32e-$ 3 when tuning $\lambda$.
Method N@10 R@10 $k$ N@10 R@10 $\lambda$ N@10 R@10
AutoSAM 0.0672 0.1117 2e-1 0.0663 0.1111 0.00 0.0663 0.1109
w/o $r^{pre}$ 0.0663 0.1109 2e-2 0.0666 0.1113 0.25 0.0668 0.1115
w/o $r^{pp}$ 0.0603 0.1016 2e-3 0.0672 0.1117 0.50 0.0672 0.1117
w/o $r$ 0.0581 0.0984 2e-4 0.0664 0.1109 0.75 0.0634 0.1057
2e-5 0.0596 0.0995 1.00 0.0603 0.1016
4.5. Ablation Study
4.5.1. Ablation Study of Sampling Reward.
In the left column of Table 4, we analyze the efficacy of each component of the reward on the Alipay dataset. From the results, we can find that removing either of the components decreases the performance. Besides, $r^{pp}$ plays a more important role than $r^{pre}$. A possible reason is that context coherence is more targeted at a single item which could more directly reflect its importance and representative. We further show the impact of scaling factor $k$ and trade-off parameter $\lambda$ by performing a grid search from {$2e^{-1}$, $2e^{-2}$, $2e^{-3}$, $2e^{-4}$,$2e^{-5}$} and $\{0,0.25,0.5,0.75,1\}$, respectively. The model achieves the best performance at $(k,\lambda)=(2e^{-3},0.5)$. The results demonstrate the effectiveness of the incorporation of these two aspects, and it also exhibits the stability of AutoSAM while the different settings of $k$ within the appropriate range bring few influence.
(a) Tmall
(b) Alipay
(c) Yelp
(d) Amazon
Figure 4. Performance and computation cost w.r.t. sample rates (portion of the kept items) by setting relax factor from $\{-0.5,0,...,2\}$. The dashed line represents the baseline model, i.e., FullSAM.
4.5.2. Efficiency and Effectiveness of Different Relax Factors
In practice, it is necessary to balance efficiency and effectiveness when adapting to online services in different scenarios. Thus, we control the relax factors $b$ in Eq. 12 to solve this problem. The performance and the computation cost of the well-trained recommender in inference w.r.t. average sample rate by setting relax factor from $\{-0.5,0,...2\}$ are shown in Figure 4. Note that the computation cost is measured with million floating-point operations (MFLOPs). We surprisingly observe that even preserving 50% interactions in inference stage can achieve the comparable performance as baseline. However, while the sample rate reaches 80%, the performance becomes to drop which probably because the tolerant sampler leads to a uniform distribution. Besides, the shortened sequences save lots computations especially with small $b$. Since the sampling can be regarded as a binary classification problem which is much simpler than recommendation and requires only one layer, we believe AutoSAM could be more efficient than using full sequence as the model gets deeper.
(a) Tmall
(b) Alipay
Figure 5. Portion of different kind behaviors of picked and dropped interactions.
Figure 6. Case study analysis in terms of sampling probabilities (P) and book titles on Amazon Dataset.
4.6. Sample Analysis
4.6.1. Sampling Quality Evaluation.
Sampling decisions play important roles to the recommend performance while directly affecting the inputs of recommendation models (Ding et al., 2020; Zhang et al., 2021). Next, we evaluate the sample qualities by comparing the portion of different kind behaviors on sampled and dropped items. Commonly, items that are purchased or added to favorites should be of more importance than clicked ones. Note that we do not use this kind of information during the training procedure. As shown in Figure 5, the proportion of purchases among picked items is higher than that among dropped items. According to this, we can find that the learned distributions are more significant to the user’s preference, which demonstrates the effectiveness of the sampler and also implies the necessity of treating behaviors non-uniformly. As it still keeps lots of clicked interactions, a possible reason is that the click items can also reflect the user’s interests to a great extent, and the factors behind that are usually very complex in the real world.
4.6.2. Case Study
We additionally show several interesting case stduies in terms of sample probabilities on Amazon dataset in this part. Given historical behaviors, possibilities of being selected by sampler are partially shown with their titles in Figure 6. The case (a) is mainly about computer science while the case (b) is mainly about cooking. As we can see, items close to the global interests are more representative, naturally should of more chance to be selected. On the contrary, items in red seem to be irrelevant, resulting in fewer opportunities to be chosen. Such visualization also implies the importance of modeling the distributions of positive user behaviors instead of treating them as uniformly positive.
5. Related Work
5.1. Sequential Recommendation
Sequential recommendation aims to predict users’ future behaviors given their historical interaction data. Early approaches mainly fuse Markov Chain and matrix factorization to capture both long- and short-term item-item transitions (Rendle et al., 2010). Latter, with the success of neural network, recurrent neural network (RNN) methods are widely conducted in sequential recommendation (Hidasi et al., 2015, 2016). Besides, convolutional-based models (Tang and Wang, 2018; Yuan et al., 2019) can also be very effective in modeling sequential behaviors. In addition, we notice that graph neural networks (Wu et al., 2019; Xu et al., 2019) have become increasingly prevalent by constructing graph structures from session sequences. And in recent years, self-attention models (Kang and McAuley, 2018; Sun et al., 2019; de Souza Pereira Moreira et al., 2021) have shown their promising strengths in the capacity of long-term dependence modeling and easy parallel computation. Furthermore, some methods (Xie et al., 2022a; Qiu et al., 2022) design contrastive learning based module to obtain better representation for sequential behaviors. Though successful, these methods focus on the model architectures and treat all behaviors as uniformly positive, may lead to the sub-optimal performance.
5.2. User Behavior Sampling
As a data-driven technology, recommendation systems have attracted a series of works that consider from the data perspective in recent years (Qin et al., 2021). To be specific, in SIM (Pi et al., 2020), a two-stage method with a general search unit (GSU) and an exact search unit is proposed to model long-term user behaviors better. Similarly, UBR4CTR (Qin et al., 2020) conducts retrieval-based method to achieve the goal. Besides, SDIM (Cao et al., 2022) proposed a sampling-based approach which samples similar items to the target by multiple hash functions. And some other works (Wang et al., 2021; Lin et al., 2023) mainly formulate a denoising task with the aims of filtering irrelevant or noise items. Despite effectiveness, these methods are mainly designed for CTR prediction, which may be not suitable to conduct the left-to-right generative loss and sample items dynamically in the SRS task. We also notice some works which attempt to capture more informative patterns between user behaviors sequence by designing advance model architectures. RETR (Yao et al., 2022) build recommender transformer for sampling the user behavior pathway. And an all-MLP based model is proposed (Zhou et al., 2022) which adopts Fourier transform with learnable filters to alleviate the influence of the noise item. Differently, in this paper, we present a general framework to learn the distributions of the raw inputs adaptively with carefully designed multi-objectives, so as to enhance the SRS with more informative and diverse training data.
6. Conclusion
In this work, we proposed a general automatic sampling framework, named AutoSAM, to non-uniformly treat historical behaviors. Specifically, a light-weighted sampler was first leveraged to adaptively explore the distribution of raw input, so that the sequential recommender systems (SRS) could be trained with more informative and diverse samples. Considering the challenges of non-differentiable actions and the necessity to introduce multiple decision factors for sampling, we further introduced a novel reinforcement learning-based method to guide the training of the sampler. We carefully designed multi-objective rewards and combined policy gradient to optimize the whole framework in an end-to-end manner. We conducted extensive experiments on four public datasets. The experimental results showed that the AutoSAM could obtain a higher performance gain by adaptively sampling informative items. We also noticed some limitations. AutoSAM conducts a transformer-based sampler, which may not be more efficient than baseline in the shallow SRS, according to Section 3.5. We believe this could be improved by adopting much lighter samplers like RNN or MLP, We hope this paper could inspire more works to be proposed from the data perspective for SRS.
Appendix A appendix
A.1. Mathematical Derivations of $r^{pre}$
According to Eq. 7, and note that $\nabla_{\phi}log\pi_{\phi}=\frac{\nabla_{\phi}\pi_{\phi}}{\pi_{\phi}}$, therefore:
(16) $\nabla_{\phi}J(\theta,\phi)$ $=\mathbb{E}_{X}\sum_{A}\nabla_{\phi}\pi_{\phi}(A|X)\mathcal{L}(\hat{X};\theta)$
$=\mathbb{E}_{X}\sum_{A}\pi_{\phi}(A|X)\nabla_{\phi}log\pi_{\phi}(A|X)\mathcal{L}(\hat{X};\theta)$
$=\mathbb{E}_{X}\mathbb{E}_{A\sim\pi_{\phi}(A|X)}\nabla_{\phi}log\pi_{\phi}(A|X)\mathcal{L}(\hat{X};\theta),$
since $log\pi_{\phi}(A|X)=log\prod\pi_{\phi}(a_{t}|X)=\sum log\pi_{\phi}(a_{t}|X)$ , we have
(17) $\nabla_{\phi}J(\theta,\phi)$ $=\mathbb{E}_{X}\mathbb{E}_{A\sim\pi_{\phi}(A|X)}\nabla_{\phi}\sum_{t=1}^{n}log\pi_{\phi}(a_{t}|X)\mathcal{L}(\hat{X};\theta)$
$=\mathbb{E}_{X}\mathbb{E}_{A\sim\pi_{\phi}(A|X)}\nabla_{\phi}\sum_{t=1}^{n}[log\pi_{\phi}(a_{t}|X)\sum_{t=1}^{n-1}\mathcal{L}_{t}(X,A;\theta)]$
$=\mathbb{E}_{X}\mathbb{E}_{A\sim\pi_{\phi}(A|X)}\nabla_{\phi}\sum_{t=1}^{n}-log\pi_{\phi}(a_{t}|X)r^{pre}.$
A.2. Experiment Details
A.2.1. Optimal Hyper-parameter
Table 5. AutoSAM’s hyper-parameter exploration.
Hyper-parameter Tuning Range Tmall Alipay Yelp Amazon
$t$ $[1.0,3.0,5.0,7.0,9.0]$ 5.0 5.0 5.0 3.0
$b$ $[-0.5,0.0,0.5,1.0,1.5,2.0]$ 1.0 2.0 1.0 0.5
$k$ $[2e^{-1},2e^{-2},2e^{-3},2e^{-4},2e^{-5}]$ $2e^{-3}$ $2e^{-3}$ $2e^{-2}$ $2e^{-3}$
$\lambda$ $[0.25,0.5,0.75]$ $0.5$ $0.5$ $0.5$ $0.5$
$\psi_{0}$ $[0.2,0.5,0.8]$ $0.8$ $0.8$ $0.8$ $0.5$
We use grid search to find the best group of AutoSAM’s hyper-parameters shown in 5. Note that we change the threshold $\psi$ according to the following schedule: $\psi=\psi_{0}+0.03(Epoch-1)$.
(a) Alipay
(b) Yelp
Figure 7. Ablation Study of temperature.
A.2.2. Ablation Study of Temperature
The temperature $\tau$ in the Softmax function of the sampler significantly influences the distribution of sample probabilities, consequently impacting the recommendation performance of the sequential recommender system. In this part, we analyze this impact by varying $\tau$ from $\{1,3,5,7,9\}$. As shown in Figure 7, we observe that AutoSAM achieves the best performance at $\tau=5$. Deviating from this value, either by choosing too large or too small $\tau$, leads to a decline in performance. A possible explanation is that smaller $\tau$ may cause convergence to local optima, while larger $\tau$ results in nearly uniform distributions.